Grade 12th passGeneral Physics

In a regular hexagon PQRSTU, Find the vector sum of PQ+PR+PS+PT+PU in terms of vector PO where O is the centre of hexagon

Nirmal
10 Years agoGrade 12th pass

1 Answer

Santosh Datla
10 Years ago
A hexagon is a six-sided figure inscribed within a circle.
The length of each side is that of the radius

Since the center is equidistant from all vertices,
and since each pair of opposing vertices lie opposite sides
of the center, P + S + Q + T + R + U = 0

Let O be the center, PQ = TS
SO = OP = -QR. Hence, TP = TS + SP = a - 2b
Likewise, UO = PQ and OP = -QR. Hence, UP = a - b